# BLISS: A Lightweight Bilevel Influence Scoring Method for Data Selection in Language Model Pretraining

### v1

@@ 7 Conclusion
In this paper, we present BLISS, a lightweight bilevel influence scoring method for data selection in language model pretraining. BLISS utilizes a proxy model, a score model, and a novel bilevel optimization framework to capture the long-term influence of data without relying on external pretrained models. Experimental results demonstrate its effectiveness in selecting data for pretraining Pythia and LLaMA models. However, current data selection methods primarily focus on language models. In future work, we plan to extend our approach to visual or multimodal models.
@@ Appendix B Implementation Details in LLaMA Experiment
There is a little difference in how we deal with the proxy model in LLaMA setting compared to Pythia setting. In addition to resetting the proxy model (LLaMA-134M) at the beginning of each round, we reset it to the initial state every 50 steps of the update of the score model. We distill the target model into the proxy model by minimizing the KL divergence for 240 steps. Then the checkpoint of the proxy model is saved as “initial" state. Since periodic resetting the proxy model ensures a close alignment between two models, we remove the KL divergence regulation term in the lower level loss function. To achieve a better lower-level solution, the proxy model executes 4 lower-level updates, each computed on a batch of 64 samples. After the score model is trained for 50 optimization steps, we reset the proxy model to the initial state.
@@ C.2 The Size of Proxy Model
We conduct experiments using two different sizes of proxy/score models (31M and 160M) for a 410M LLM. We observe that the KL divergence between the proxy and the LLM remains low for both sizes-0.15 for the 160M model and 0.10 for the 31M model. The corresponding learning curves are shown in Figure 7(a), which presents the results from round 2. The performance comparison of two sizes of proxy model is summarized in Table 6. These findings suggest that even a small proxy model (31M) is sufficient to serve as an effective surrogate for the 410M LLM.
(a) The evolution of the lower-level training loss and KL divergence for different proxy model size. Subfigures (a), (b): Proxy model size 31M, target LLM size 410M. Subfigures (c), (d): Proxy model size 160M, target LLM size 410M.
Table 6: Comparison of BLISS with different size of proxy/score model and on zero-shot evaluation over multiuple downstream datasets (410M model, 10B tokens) with 20k-step training.
@@ Appendix D Additional results
Table 8: Results of Different Methods under the 410M/1B Setting. Subscripts denote standard deviations. Best scores are in bold.
@@ Appendix E Experimental Hyperparameters
Table 9: Experimental settings
@@
batch size 16(Pythia-410M, LLaMA-0.5B)/32(Pythia 1B)
Proxy/Score model learning rate($\eta_{1}/\eta_{2}$) 1e-5
GDLS learning rate($\eta$) 1e-2
@@ Appendix F Evolution of training and validation loss
In Figure F, 11(a), the training loss and validation loss correspond to the lower-level and upper-level objective functions in Equation (1), respectively.
(a) The evolution of the lower-level training loss and upper-level validation loss in round 2.
(a) The evolution of Lower-level training loss and upper-level validation loss in round 5.
(b) The visualization of lower-level training loss and the upper-level validation loss in round 2 bilevel optimization without KL divergence.
@@ Appendix G Distributed Softmax to Compute Influence Score
$P_{i}=\frac{e^{h(\theta_{s};\xi_{i})}}{\sum_{j=1}^{B}e^{h(\theta_{s};\xi_{j})}}$ (7)

### v3

@@ 7 Conclusion
In this paper, we present BLISS, a lightweight bilevel influence scoring method for data selection in language model pretraining. BLISS utilizes a proxy model, a score model, and a novel bilevel optimization framework to capture the long-term influence of data without relying on external pretrained models. Experimental results demonstrate its effectiveness in selecting data for pretraining Pythia and LLaMA models. However, the bilevel optimization in our method may incur higher peak memory usage, which remains an important direction for future work.
@@ Appendix B Implementation Details in LLaMA Experiment
There is a little difference in how we deal with the proxy model in LLaMA setting compared to Pythia setting. In addition to resetting the proxy model (LLaMA-134M) at the beginning of each round, we reset it to the initial state every 50 steps of the update of the score model. We distill the target model into the proxy model by minimizing the KL divergence for 240 steps. Then the checkpoint of the proxy model is saved as “initial" state. Since periodically resetting the proxy model ensures a close alignment between two models, we remove the KL divergence regularization term in the lower level loss function. To achieve a better lower-level solution, the proxy model executes 4 lower-level updates, each computed on a batch of 64 samples. After the score model is trained for 50 optimization steps, we reset the proxy model to the initial state.
@@ Appendix C More Ablation Studies
Table 5: Comparison of BLISS with different settings(without softmax and single level update) over multiple downstream datasets (410M model, 10B tokens) with 20k-step training.
@@ C.2 The Size of Proxy Model
We conduct experiments using two different sizes of proxy/score models (31M and 160M) for a 410M LLM. We observe that the KL divergence between the proxy and the LLM remains low for both sizes-0.15 for the 160M model and 0.10 for the 31M model. The corresponding learning curves are shown in Figure 4, which presents the results from round 2. The performance comparison of two sizes of proxy model is summarized in Table 6. These findings suggest that even a small proxy model (31M) is sufficient to serve as an effective surrogate for the 410M LLM.
(a) Proxy size 31M, target LLM 410M.
(b) Proxy size 160M, target LLM 410M.
Figure 4: Evolution of lower-level training loss and KL divergence for different proxy model sizes.
Table 6: Comparison of BLISS with different size of proxy/score model and on zero-shot evaluation over multiple downstream datasets (410M model, 10B tokens) with 20k-step training.
@@ Appendix D Additional results
Table 8: Results of Different Methods under the 410M/1B Setting. Subscripts denote standard error. Best scores are in bold.
@@ Appendix E Experimental Hyperparameters
Table 9: Experimental settings.
@@
batch size 16(Pythia-410M)/32(Pythia 1B)
Proxy/Score model learning rate($\eta_{1}/\eta_{3}$) 1e-5
GDLS learning rate($\eta_{2}$) 1e-2
@@ Appendix F Evolution of training and validation loss
In Figure 6(a), 6(b), we visualize the curves training loss pretraining round 2 and 5.
(a) Round 2.
(b) Round 5.
Figure 6: Evolution of training loss within a bilevel training round.
(a) Training loss vs. steps.
(b) Validation loss vs. steps.
Figure 7: Lower-level training loss and upper-level validation loss in round 2 of bilevel optimization without KL divergence.
@@ Appendix G Distributed Softmax to Compute Influence Score
$P_{i}=\frac{e^{h(\theta_{s};\xi_{i})}}{\sum_{j=1}^{B}e^{h(\theta_{s};\xi_{j})}}.$ (6)
@@
Appendix H Running Time and Memory
We measured the memory and runtime of the data selection stage for both BLISS and MATES under different target (or pretraining) model sizes (for short, T: target). The results are shown in Table˜10. We have two observations:
  • BLISS scales well with larger target models. Note that the target model is not an optimization variable for the bilevel optimization and it is only used for calculating the KL divergence. Therefore, it does not affect the scalability of bilevel optimization. When increasing the target from 410M to 1B, BLISS’s memory and runtime grow moderately ($49.46\rightarrow 74.51$ GB; $5.03\rightarrow 11.82$ hours), as expected.
  • BLISS is significantly faster than MATES. MATES incurs high cost because each round requires oracle data collection. For every example, MATES performs a one-step gradient update on the target model and evaluates the validation loss change to compute influence scores. This per-example simulation dominates runtime. In contrast, BLISS avoids all per-example oracle evaluations in MATES and performs bilevel optimization directly on the proxy/score model, leading to $2-3\times$ faster data selection.
Table 10: The comparison of runtime and memory in data selection stage.
  Setting Memory peak (GB) Total data selection time (hours)
MATES: T: 410M $36.36$ $18.0$
MATES: T: 1B $63.52$ $30.32$
BLISS: T: 410M $49.46$ $5.03$
BLISS: T: 1B $74.51$ $11.82$
Appendix I Ablation Study for Bilevel Optimization Steps
We did the ablation study to investigate the steps of bilevel optimization on Pythia models. As shown in Figure˜9, both the 3k-step and 5k-step settings converge to nearly the same training loss. This indicates that 3k steps are sufficient for the Pythia-type proxy model, as increasing the steps to 5k does not yield additional improvements. So we fix the training steps of bilevel optimization to 3k steps in main experiments.
Figure 8: Training loss with different steps.
Figure 9: Test loss on SlimPajama-6B.
Appendix J Domain Reweighting: Fidelity of Proxy Models
To verify the fidelity of proxy models to full-scale LLMs, we conduct a domain-reweighting experiment on the SlimPajama-6B dataset (DKYoon, 2023), which is a multi-domain dataset, including ArXiv, Books, C4, CommonCrawl, GitHub, StackExchange, and Wikipedia. The objective is to learn optimal domain weights $\alpha\in\mathbb{R}^{d}$, $d$ is the domain size, such that a model trained on data sampled according to the weights achieves the best downstream performance.
We compare two settings:
  1. 1.
    Case 1 (with proxy model): The lower level optimizes a lightweight proxy model (LLaMA-134M) with output alignment to the target LLM (LLaMA-300M), and the upper level learns the domain weights $\tilde{\alpha}$.
  2. 2.
    Case 2 (without proxy model): The lower level directly optimizes the target LLM (LLaMA-300M), and the upper level learns the domain weights $\alpha$.
We perform bilevel optimization for 1,000 steps in both cases to learn the domain weights, where $10\%$ of the original training set is held out as the validation set for the upper-level objective, and the remaining $90\%$ is used as the lower-level training set. After obtaining $\tilde{\alpha}$ and $\alpha$, we train two final LLaMA-300M models on data sampled according to each set of weights, respectively . Figure 10 presents the learning curves of domain weights for both cases. We observe that the trajectories of $\tilde{\alpha}$ and $\alpha$ are highly similar across most domains (e.g., in the domain of Wikipedia, Book, Stackexchange), demonstrating that the proxy model maintains high fidelity to the full-scale LLM in data selection.
Finally, we evaluate the resulting pretrained LLMs on the test set of SlimPajama-6B, and the results are shown in Figure 9. The test loss curves show that data selection based on the proxy model maintains high fidelity to the full-scale LLM, while significantly outperforming random selection.
(a) Wikipedia
(b) Book
(c) StackExchange
(d) CommonCrawl
(e) ArXiv
(f) GitHub
Figure 10: Learning curves of domain weights.